OpenStudy (anonymous):

Find the slope of the line passing through (11, -2) and (2, -2). A. 4/9 B. 9 C. 0 D. Undefined

OpenStudy (jack1):

y = mx + c m = slope = y2 - y1 / x2 - x1

OpenStudy (jack1):

so your points on the graph are labeled as ( x, y) so your first point is 11(x1), -2 (y1) and your 2nd is 2 (x2), -2 (y2)

OpenStudy (jack1):

so what's your slope equal?

OpenStudy (jack1):

@Zoobidoo

OpenStudy (jack1):

ooo...shiny, thanks ;)

OpenStudy (anonymous):

Hello sorry for taking so long. I logged out.

OpenStudy (jack1):

sçool

OpenStudy (anonymous):

Im working on it.

OpenStudy (anonymous):

would it be 9?

OpenStudy (jack1):

no dude, sorry y2 -y1 / x2 - x1 =-2 - (-2) / 2- 11 = 0 / -9 = 0

OpenStudy (anonymous):

oh ok ok, I see.

OpenStudy (jack1):



OpenStudy (jack1):

cool, catchya

OpenStudy (anonymous):

ok, yes I was completely off.

OpenStudy (anonymous):

Thank you for your help @Jack1

OpenStudy (jack1):

all good dude, always welcome